OpenStudy (anonymous):

(sin2x)^2-4sin2x-1=0 Will add all the steps I took

OpenStudy (anonymous):

\[(\sin2x)^{2}-4\sin2x=1\] \[(\sin2x)^{2}-4\sin2x=1\] +4 +4 \[\sqrt{(\sin2x)^{2}-\sin2x}==\sqrt{5}\] \[(\sin2x)-2=+-\sqrt{5}\] +2 +2 \[(\sin2x)=2+-\sqrt{5}\] \[2x=(sin)^{-1}(2+-\sqrt{5})\] Now wut?..